OpenStudy (anonymous):

Write the first five terms of the arithmetic sequence with the first term a1 and common difference d.

OpenStudy (anonymous):

a1=1/13 d =3/13

OpenStudy (mertsj):

Do you understand that for an arithmetic sequence you start with the first term and then add d to get the second term?